OpenStudy (arisa.s):

Need help with a question related to matrices. Our instructor didn't teach us how to do this but it's in our assignment as a compulsory question. Let M= [-2 1, -36 10]. Find formulas for M^n where n is a positive integer. M^n=[? ?, ? ?]

OpenStudy (arisa.s):

Here's the question as it's shown in the assignment.

OpenStudy (mathmale):

Assume that we want to find \[M ^{1}, M^2, M^3, ...\]

OpenStudy (mathmale):

... where

OpenStudy (mathmale):

the first is just the matrix M itself, and \[M^2\] ... is the square of matrix M, and so on.

OpenStudy (mathmale):

Take matrix M and multiply it by itself, of course following the rules for matrix multiplication. Try this, please, showing all work on which you'd like to receive feedbacl/

OpenStudy (arisa.s):

Okay, I got M^2 as [-32 8, -288 64]. Now what should I do?

OpenStudy (mathmale):

I'm going to assume that your answer is correct. To find M^3, multiply your most recent matrix result by M.

OpenStudy (arisa.s):

M^3=[-224 48, 1728 352]

OpenStudy (mathmale):

Glad to say, we've both gotten the same result for M^2. I won't be doing the entire problem myself, but encourage you to find M^3 and M^4. Your next step is to look at the element in the first row, first column of M^1, M^2, M^3 and M^4 to determine whether that element seems to be predictable depending upon the exponent n.

OpenStudy (mathmale):

If you have a TI-83 or -84 or more advanced, more recent calculator, you can do matrix multiplication with it. Strongly encourage you to learn how to do this.

OpenStudy (mathmale):

Notice that this problem asks you for the contents of four different "answer boxes." This means you need to try to find a formula for each such box, so that you can predict the numerical value of any such box if given 2x2 matrix M and exponent n.

OpenStudy (unklerhaukus):

\[\textbf M= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\] \[\textbf M^2=\textbf M\textbf M= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\] \[\textbf M^3=\textbf M\textbf M\textbf M= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\]

OpenStudy (unklerhaukus):

you have \(\textbf M ^2\) correct

OpenStudy (unklerhaukus):

\[\textbf M^2=\textbf M\textbf M= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix} =\begin{bmatrix}-32 &8\\ -288& 64\end{bmatrix}\] \[\textbf M^3=\textbf M\textbf M^2= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\begin{bmatrix}-32 &8\\ -288& 64\end{bmatrix}\]

OpenStudy (mathmale):

Unkle: Your input is helpful and surely looks good. All I ask is that you not do for the other person what that person could do for himself or herself. Thanks.

OpenStudy (unklerhaukus):

Arisa.S already has shown the correct answers for M^2,^3

OpenStudy (unklerhaukus):

except there is a sign error in M^3

OpenStudy (mathmale):

Unkle: Thanks for noticing that. Perhaps you could engage Arisa in locating and fixing the sign error.

OpenStudy (arisa.s):

it's supposed to be -1728 in the second row first column, right?

OpenStudy (unklerhaukus):

that's it

OpenStudy (unklerhaukus):

So you have \[\textbf M= \begin{bmatrix}-2 &1\\ -36 &10\end{bmatrix}\]\[\textbf M^2=\begin{bmatrix}-32 &8\\ -288& 64\end{bmatrix}\]\[\textbf M^3=\begin{bmatrix}-244&48\\-1728&352\end{bmatrix}\]

OpenStudy (unklerhaukus):

hmm, this doesn't resolve into a nice pattern, i think this is a dead end. We might need a different method

OpenStudy (unklerhaukus):

Can you diagonalize the matrix \(\mathbf M\)?

OpenStudy (mathmale):

Agreed...no distinct pattern has hit me yet. Pls explain to Arisa what you mean by "diagonalize the matrix M," and again to join our dialogue.

OpenStudy (arisa.s):

I couldn't figure out a relation either.

OpenStudy (thomas5267):

Diagonalisation is the process of finding a matrix \(\mathbf{P}\) such that \[ \mathbf{A}=\mathbf{PD}\mathbf{P}^{-1}\text{ or }\mathbf{D}=\mathbf{P^{-1}}\mathbf{AP} \] \(\mathbf{D}\) of course is a diagonal matrix. If \(\mathbf{A}=\mathbf{PD}\mathbf{P}^{-1}\), then \(\mathbf{A}^n=\mathbf{PD}^n\mathbf{P^{-1}}\). The only problem is that this matrix is not diagonalisable. Have your teacher taught you the Jordan normal form yet?

OpenStudy (thomas5267):

I used the computer software Mathematica to generate a closed-form solution but I don't see how you can get it without using Jordan normal form.